OpenStudy (anonymous):

Jarred sells DVDs. His inventory shows that he has a total of 3,500 DVDs. He has 2,342 more contemporary titles than classic titles. Let x represent the number of contemporary titles and y represent the number of classic titles. The system of equations models the given information for both types of DVDs. x + y = 3,500 x – y = 2,342 Solve the system of equations. How many contemporary titles does Jarred have? A. 1,158 B. 1,737 C. 2,342 D. 2,921

OpenStudy (retireed):

Add the two equations. The y will be eliminated.

OpenStudy (retireed):

Can you do this ... x + y = 3,500 + x – y = 2,342 ----------------

OpenStudy (retireed):

Still waiting ....

OpenStudy (anonymous):

I got it, it's 2,921 thank you!

OpenStudy (retireed):

Very good!